OpenStudy (anonymous):

62.5% is equivalent to what fraction in reduced terms?

OpenStudy (dumbcow):

when converting percent to fraction, divide by 100 62.5/100 multiply by 10/10 to get rid of decimal =625/1000 Reduce by cancelling out common factors 625 = 25*25 = 5*5*5*5 1000 = 10*10*10 = 5*5*5*2*2*2 3 5's cancel on top and bottom = 5/8